OpenStudy (anonymous):

If k and h are functions, then the composition function k \circle h and the composition function h \circle k are ________ .

OpenStudy (anonymous):

never equal to each other always equal to each other equal to each other only when k(x) = h(x) sometimes equal to each other

OpenStudy (anonymous):

I say equal to eachother only when k(x)= h(x)

OpenStudy (goformit100):

equal to each other only when k(x) = h(x) ....................is correct

OpenStudy (goformit100):

done